OpenStudy (anonymous):

Log(1-2x) = 1 Find x.

OpenStudy (anonymous):

Whenever they don't tell you what the base of the log is, just assume it's 10. Log(1-2x) = 1 10^1 = 1-2x 9=-2x x=-9/2 To confirm that this is correct, put it back into the original equation and see if you get 1 Log(1-2(-9/2) = Log(1+9) = Log(10) Log(10) = 1 Which is true.